# 3D Shape Completion on Unseen Categories: A Weakly-supervised Approach

### v1

@@ Abstract
3D shapes captured by scanning devices are often incomplete due to occlusion. 3D shape completion methods have been explored to tackle this limitation. However, most of these methods are only trained and tested on a subset of categories, resulting in poor generalization to unseen categories. In this paper, we introduce a novel weakly-supervised framework to reconstruct the complete shapes from unseen categories. We first propose an end-to-end prior-assisted shape learning network that leverages data from the seen categories to infer a coarse shape. Specifically, we construct a prior bank consisting of representative shapes from the seen categories. Then, we design a multi-scale pattern correlation module for learning the complete shape of the input by analyzing the correlation between local patterns within the input and the priors at various scales. In addition, we propose a self-supervised shape refinement model to further refine the coarse shape. Considering the shape variability of 3D objects across categories, we construct a category-specific prior bank to facilitate shape refinement. Then, we devise a voxel-based partial matching loss and leverage the partial scans to drive the refinement process. Extensive experimental results show that our approach is superior to state-of-the-art methods by a large margin. We will make the source code publicly available at https://github.com/ltwu6/WSSC.
@@ I Introduction
In recent years, with the popularity of 3D scanning equipment, 3D sensing technology has developed rapidly. 3D shapes, such as point clouds, meshes and 3D voxels, have become increasingly valuable in various downstream tasks, such as SLAM, virtual reality, and autonomous driving [1, 2, 3, 4, 5, 6, 7]. However, obtaining high-quality 3D shapes directly from 3D scanning devices is challenging in practical applications. Due to self-occlusion as well as occlusion from environmental obstacles, the captured raw 3D data typically suffer from incompleteness, which may significantly impact downstream task performances.
@@
In practical scenarios, however, many categories of partial scans lack complete ground truth due to limited resources such as human labor and time. Consequently, supervised shape completion methods exhibit poor generalization to unseen categories. To solve this problem, there have emerged some unsupervised shape completion methods [17, 18, 19, 20, 21, 22]. These unsupervised shape completion methods can be divided into two categories according to their training data. The first category employs data from the target categories for both training and testing, while the second category uses data from the same category but different domains for training. The former ignores shape prior information from the existing data categories, while the latter only focuses on completing data from different domains rather than different categories. Given partial-complete pairs from a set of categories, we argue that leveraging their prior knowledge can enable the completion of partial shapes from novel categories.
@@
In this paper, we propose a weakly-supervised method to infer complete shapes from unseen categories. We first devise an end-to-end Prior-assisted Shape Learning Network (PSLN) to predict a coarse complete shape in a supervised manner. Specifically, we construct a prior bank composed of some representative complete shapes from the seen categories. Next, we design a Multi-scale Pattern Correlation (MPC) module to capture multi-scale local patterns associated with the input shape from the priors. Subsequently, we leverage these learned local patterns to effectively reconstruct the complete shape of the partial input. Different from PatchComplete, our PSLN computes the correlation between the priors and input at the feature level instead of in Euclidean space. This design enables our model to be trained end-to-end in a single pass. Furthermore, since the input and priors may exhibit similar local patterns at varying scales, such as the resemblance between the large lampstand of some streetlights and the small legs of chairs, we extract priors’ local patterns of different scales and compute the correlations between them and partial input at each layer of MPC. Experimental results demonstrate that our PLSN achieves superior performance compared to PatchComplete while significantly reducing computational and storage requirements.
Moreover, inspired by self-supervised point cloud completion methods [23, 24, 25], we leverage the information of partial scans from unseen categories to refine the coarse outputs. These self-supervised point cloud methods use the partial scans of the input object along with the partial matching loss to guide the model to predict the complete shape. However, as the loss functions for self-supervised point cloud completion are unsuitable for voxel grids, we propose a voxel-based partial matching loss to guide the refinement process. Additionally, we believe that objects within a category exhibit more similar local patterns than those across different categories. Therefore, we design specific priors for each category during the refinement stage. We refer to this refinement framework as Category-specific Shape Refinement Network (CSRN). Our approach utilizes complete shapes from seen categories and partial scans from unseen categories as supervision. Since this supervisory information is indirect and partial in guiding the complete shapes from unseen categories, our approach can be regarded as a form of weakly-supervised method.
@@
  • We propose a Category-specific Shape Refinement Network that refines the coarse shapes with partial scans from unseen categories in a self-supervised manner. This model takes into account the shape characteristics of different categories and accurately performs category-specific refinement for the coarse shapes.
The remainder of this paper is organized as follows. Section II reviews the previous supervised and unsupervised shape completion and reconstruction approaches. Section III introduces our methods in detail. Section IV introduces our experimental settings, results and analyses. Section V concludes this work.
@@ II-A Supervised 3D Shape Completion and Reconstruction
Voxel grids can be seen as an extension of the 2D pixel representation to 3D image grids. It is widely used due to its simple and regular representation. The grids are filled with occupancy values or distance field values. The occupancy value of a voxel is typically binary, with a value of 0 representing an empty voxel and a value of 1 indicating an occupied voxel. In certain cases, the occupancy value can be represented as a continuous probability distribution, taking on values between 0 and 1, which reflect the probability that the voxel is occupied. Tatarchenko et al. [28] introduced a novel deep neural network architecture, termed Octree Generating Network (OGN), which can efficiently generate high-resolution 3D outputs via an octree-based representation. By utilizing hierarchical octree structures to represent 3D data, OGN is able to perform sparse computations and process large-scale 3D data efficiently. Sun et al. [26] proposed to exploit the local patches of partial inputs to infer the missing parts, since the local patterns of an object may repeat. They first use a network to retrieve the most correlated local patches from inputs and then deform and blend them into the final outputs.
@@
Distance field values encode the distance from a point in space to the closest surface of an object. Dai et al. [27] utilized a 3D-Encoder Predictor Network to generate an initial low-resolution shape prediction. Subsequently, a 3D shape synthesis method was devised to refine the preliminary output by integrating shape priors obtained from a complete shape database. Chibane et al. [29] proposed a 3D shape reconstruction and completion model named Implicit Feature Networks (IF-Nets), which can consume various 3D inputs, preserve objects’ details and reconstruct articulated humans. They use multi-scale 3D vectors to encode 3D shapes and predict the occupancy at the feature space. Mittal et al. [8] introduced AutoSDF, a method that leverages an autoencoder to acquire a compact latent representation of 3D shapes, which in turn generates a Signed Distance Function (SDF) that encapsulates shape priors. The generated SDF is subsequently employed to steer the 3D completion, reconstruction, and generation process. Rao et al. [9] proposed that although 3D objects may differ in global shapes, they could share similar local structures across different categories. To address this, the authors proposed a framework named PatchComplete that learns patch-based local priors from training categories and leverages these priors to infer complete shapes from unseen categories. Nevertheless, on the one hand, the training process of PatchComplete is time-consuming. On the other hand, PatchComplete used a unified seen-category prior for all categories of data, which limited its generation ability.
@@
Continuous implicit functions represent objects as mathematical functions, such as neural network models, that implicitly describe the object’s surface as the zero level-set of the function. Park et al. [46] introduced a technique for learning continuous signed distance functions (SDFs) to represent 3D shapes by training a deep neural network to map 3D points to a signed distance value. Instead of representing each object with a single neural network and a global latent code, Chabra et al. [45] proposed to split the shape’s surface into local regions and employ independent local latent codes to represent them. To solve the imbalanced distributions between different local regions, Wang et al. [13] proposed to represent an object with clusters of local regions along with multiple decoders, where each decoder is responsible for one cluster with the same pattern.
Figure 1: Flowchart of the Prior-assisted Shape Learning Network (PSLN). “Conv” and “De-Conv” denote the convolution and de-convolution operation, respectively. The PSLN leverages a Multi-scale Pattern Correlation (MPC) module to capture the associated patterns for each local region of the partial input from the priors and subsequently use these patterns to infer the complete shape hierarchically. The MPC consists of two encoders and three Cross-Attention modules. The prior encoder utilizes multiple kernels of varying sizes to capture local structures at different scales. The Cross-Attention module identifies associated local patterns by calculating attention scores between the multi-scale local features of the partial input and the priors.
@@ II-B Unsupervised 3D Shape Completion and Reconstruction
The in-domain methods [17, 18, 19, 25, 24, 23] do not use partial-complete pairs from other domains, and the data categories for the training and test set are consistent. In this paper, we mainly introduce the self-supervised point cloud methods that we have used. Mittal et al. [23] proposed the first self-supervised point cloud completion method, PointPnCNet. PointPnCNet learns to infer complete shapes from partial scans with complete ground-truth shapes as supervision. It randomly removes a region from the partial scan, employs neural networks to predict the missing part, and leverages partial matching loss and partial scans as supervision. On the one hand, the network tends to extract common patterns among shapes in the same category. On the other hand, the model is unable to distinguish between synthetic removal and natural incompleteness. As a result, PointPnCNet is capable of successfully predicting complete shapes. Wang et al. [24] proposed a two-branch cascaded refinement network for point cloud completion. In order to learn to predict complete shapes in the absence of ground truth, they remove a portion of the area from partial inputs and combine them with the MixUp [51] strategy as inputs and supervisions. Gu et al. [25] proposed a weakly-supervised model that can simultaneously predict complete shapes and 6-DoF poses with the help of multi-view partial scans. Nonetheless, the performances of these in-domain unsupervised methods are constrained as they only utilize a limited number of partial shapes from fixed categories for training.
@@ III Proposed Method
We represent the shapes as 3D voxel grids.
Denote by $\mathbf{S}_{s}\in\mathbb{R}^{N\times N\times N}$ the partial shape of an object from the seen categories and $\mathbf{G}\in\mathbb{R}^{N\times N\times N}$ the corresponding ground truth, $\mathbf{S}_{u}\in\mathbb{R}^{N\times N\times N}$ the partial shape from unseen categories, where $N$ represents the resolution of the 3D voxels. Our goal is to find a mapping $f(\mathbf{S}_{u}|\mathbf{S}_{s},G)=\mathbf{G}$.
Our approach consists of two models, i.e., a Prior-assisted Shape Learning Network (PSLN) and a Category-specific Shape Refinement Network (CSRN).
III-A Multi-scale Prior-assisted Shape Learning
In real life, many 3D objects have similar local patterns. For example, most desks, benches, and lamps have sticks as support. These local structures are important for understanding the shapes of objects from unseen categories and can be used to infer missing parts of these objects. Inspired by this, we first construct a prior bank ${\mathbf{B}}$ that consists of several representative complete shapes from training data. Following PatchComplete, we use the mean-shift clustering algorithm to select $M$ complete shapes from training data as priors. These priors contain typical local structures from each category in the training data.
Next, we design an end-to-end network to learn the local patterns of the partial shape from the priors. As depicted in Figure 1, we start by using two encoders to extract local features of the partial input and priors. The features of partial input and priors at the $i$-th layer are denoted by $\mathbf{Q}_{i}\in\mathbb{R}^{N_{i}\times N_{i}\times N_{i}\times C_{i}}$ and $\mathbf{K}_{i,m}\in\mathbb{R}^{N_{i}\times N_{i}\times N_{i}\times C_{i}}$, respectively, where $m=1,2,...,M$. Then, we use cross-attention [52] to compute the weights between $\mathbf{Q}_{i}$ and $\mathbf{K}_{i,m}$ to represent their correlation. Specifically, we first split $\mathbf{Q}_{i}$ and $\mathbf{K}_{i,m}$ into vectors of size $1\times 1\times 1\times C_{i}$. We define the vectors of $\mathbf{Q}_{i}$ as ”queries” and the vectors of $\mathbf{K}_{i,m}$ as ”keys” and ”values”. Then we compute the dot products of each query with all keys and divide them by $C_{i}/2$. Next, we adopt a Softmax function to obtain the weights.
$w_{i,m}=\texttt{Softmax}\left(\frac{\mathbf{Q}_{i}\cdot\mathbf{K}_{i,m}}{d/2}\right)\,.$ (1)
Finally, we assemble the prior embeddings by:
@@ III-A Multi-scale Prior-assisted Shape Learning
where $\mathbf{V}$ represents the ”values” and $\mathbf{V}$ is equal to $\mathbf{K}$ in practice.
Additionally, we observe that local patterns of partial shapes and priors may be similar at different scales. For example, the sticks of a chair backrest may have a shape similar to that of the pole of a street lamp, but their sizes are significantly different. To address this issue, we use convolutional kernels of various sizes to extract multi-scale features of priors and then calculate the attention between all multi-scale prior features and the features of the partial input. At the first layer of the prior encoder, we apply convolutional kernels of sizes $7\times 7\times 7$, $5\times 5\times 5$ and $3\times 3\times 3$, respectively. At the second layer, the sizes of convolutional kernels are $5\times 5\times 5$ and $3\times 3\times 3$. At the last layer, we only use a single convolutional kernel of size $3\times 3\times 3$ due to the large receptive field of this layer. The multi-scale features of each layer are concatenated and passed to the next layer. We call these blocks Multi-scale Structure Learning unit (MSL).
The assembled prior embeddings $\mathbf{F}_{i}$ are then decoded into the complete shape hierarchically. Specifically, we first directly decode $\mathbf{F}_{3}$ by a deconvolution layer. Then $\mathbf{F}_{2}$ and $\mathbf{F}_{1}$ are respectively concatenated with the decoded features and processed by two deconvolution layers. The model outputs the coarse complete shape after one more deconvolution layer and a Softmax layer. The training process of the PSLN is driven by optimizing L1 loss between the reconstructed coarse shape and its corresponding ground-truth complete shape.

### v2

@@ Abstract
3D shapes captured by scanning devices are often incomplete due to occlusion. 3D shape completion methods have been explored to tackle this limitation. However, most of these methods are only trained and tested on a subset of categories, resulting in poor generalization to unseen categories. In this paper, we propose a novel weakly-supervised framework to reconstruct the complete shapes from unseen categories. We first propose an end-to-end prior-assisted shape learning network that leverages data from the seen categories to infer a coarse shape. Specifically, we construct a prior bank consisting of representative shapes from the seen categories. Then, we design a multi-scale pattern correlation module for learning the complete shape of the input by analyzing the correlation between local patterns within the input and the priors at various scales. In addition, we propose a self-supervised shape refinement model to further refine the coarse shape. Considering the shape variability of 3D objects across categories, we construct a category-specific prior bank to facilitate shape refinement. Then, we devise a voxel-based partial matching loss and leverage the partial scans to drive the refinement process. Extensive experimental results show that our approach is superior to state-of-the-art methods by a large margin. We will make the source code publicly available at https://github.com/ltwu6/WSSC.
@@ I Introduction
In recent years, with the popularity of 3D scanning equipment, 3D sensing technology has developed rapidly. 3D shapes, such as point clouds, meshes, and 3D voxels, have become increasingly valuable in various downstream tasks, such as SLAM, virtual reality, and autonomous driving [1, 2, 3, 4, 5, 6, 7]. However, obtaining high-quality 3D shapes directly from 3D scanning devices is challenging in practical applications. Due to self-occlusion as well as occlusion from environmental obstacles, the captured raw 3D data typically suffer from incompleteness, which may significantly impact downstream task performances.
@@
In practical scenarios, however, many categories of partial scans lack complete ground truth due to limited resources such as human labor and time. Consequently, supervised shape completion methods exhibit poor generalization to unseen categories. To solve this problem, there have emerged some unsupervised shape completion methods [17, 18, 19, 20, 21, 22]. These unsupervised shape completion methods can be divided into two types according to their training data. The first type employs data from the target categories for both training and testing, while the second type uses data from the same category but different domains for training. The former ignores shape prior information from the existing data categories, while the latter only focuses on completing data from different domains rather than different categories. Given partial-complete pairs from a set of categories, we argue that leveraging their prior knowledge can enable the completion of partial shapes from novel categories.
@@
In this paper, we propose a weakly-supervised method to infer complete shapes from unseen categories. We first devise an end-to-end Prior-assisted Shape Learning Network (PSLN) to predict a coarse complete shape in a supervised manner. Specifically, we construct a prior bank composed of some representative complete shapes from the seen categories in the training data. The PSLN consumes as input a partial shape along with all the priors stored in this bank. Next, we design a Multi-scale Pattern Correlation (MPC) module to capture multi-scale local patterns associated with the input shape from the priors. Subsequently, we leverage these learned local patterns to effectively reconstruct the complete shape of the partial input. Different from PatchComplete, our PSLN computes the correlation between the priors and partial shape at the feature level instead of in Euclidean space. This design enables our model to be trained end-to-end in a single pass. Furthermore, since the input and priors may exhibit similar local patterns at varying scales, such as the resemblance between the large lampstand of some streetlights and the small legs of chairs, we extract priors’ local patterns of different scales and compute the correlations between them and partial input at each layer of MPC. Experimental results demonstrate that our PSLN achieves superior performance compared to PatchComplete while significantly reducing computational and storage requirements.
Moreover, inspired by self-supervised point cloud completion methods [23, 24, 25], we leverage the information of partial scans from unseen categories to refine the coarse outputs. These self-supervised point cloud methods use the partial scans of the input object along with the partial matching loss to guide the model in predicting the complete shape. However, as the loss functions for self-supervised point cloud completion are unsuitable for voxel grids, we propose a voxel-based partial matching loss to guide the refinement process. Additionally, we believe that objects within a category exhibit more similar local patterns than those across different categories. Therefore, we design specific priors for each category during the refinement stage.
Our approach utilizes complete shapes from seen categories and partial scans from unseen categories as supervision. Since this supervisory information is indirect and partial in guiding the complete shapes from unseen categories, our approach can be regarded as a form of weakly-supervised method.
@@
  • We propose a Category-specific Shape Refinement approach that refines the coarse shapes with partial scans from unseen categories in a self-supervised manner. This model takes into account the shape characteristics of different categories and accurately performs category-specific refinement for the coarse shapes.
The remainder of this paper is organized as follows. Section II reviews the previous supervised and unsupervised shape completion and reconstruction approaches. Section III introduces our methods in detail. Section IV introduces our experimental settings, results, and analyses. Section V concludes this work.
@@ II-A Supervised 3D Shape Completion and Reconstruction
Voxel grids can be seen as an extension of the 2D-pixel representation to 3D image grids. It is widely used due to its simple and regular representation. The grids are filled with occupancy values or distance field values. The occupancy value of a voxel is typically binary, with a value of 0 representing an empty voxel and a value of 1 indicating an occupied voxel. In certain cases, the occupancy value can be represented as a continuous probability distribution, taking on values between 0 and 1, which reflect the probability that the voxel is occupied. Tatarchenko et al. [28] introduced a novel deep neural network architecture, termed Octree Generating Network (OGN), which can efficiently generate high-resolution 3D outputs via an octree-based representation. By utilizing hierarchical octree structures to represent 3D data, OGN is able to perform sparse computations and process large-scale 3D data efficiently. Sun et al. [26] proposed to exploit the local patches of partial inputs to infer the missing parts since the local patterns of an object may repeat. They first use a network to retrieve the most correlated local patches from inputs and then deform and blend them into the final outputs. Similar to [26], we also use a binary voxel grid to represent 3D objects. However, unlike existing voxel-based methods, our approach focuses on completing partial scans of unseen categories. Moreover, we leverage priors from the training set to learn the shapes from unseen categories in a weakly-supervised manner.
@@
Distance field values encode the distance from a point in space to the closest surface of an object. Dai et al. [27] utilized a 3D-Encoder Predictor Network to generate an initial low-resolution shape prediction. Subsequently, a 3D shape synthesis method was devised to refine the preliminary output by integrating shape priors obtained from a complete shape database. Chibane et al. [29] proposed a 3D shape reconstruction and completion model named Implicit Feature Networks (IF-Nets), which can consume various 3D inputs, preserve objects’ details and reconstruct articulated humans. They use multi-scale 3D vectors to encode 3D shapes and predict the occupancy at the feature space. Park et al. [46] introduced a technique for learning continuous signed distance functions (SDFs) to represent 3D shapes by training a deep neural network to map 3D points to a signed distance value. Mittal et al. [8] introduced AutoSDF, a method that leverages an autoencoder to acquire a compact latent representation of 3D shapes, which in turn generates a Signed Distance Function (SDF) that encapsulates shape priors. The generated SDF is subsequently employed to steer the 3D completion, reconstruction, and generation process. Rao et al. [9] proposed that although 3D objects may differ in global shapes, they could share similar local structures across different categories. To address this, the authors proposed a framework named PatchComplete that learns patch-based local priors from training categories and leverages these priors to infer complete shapes from unseen categories. Nevertheless, on the one hand, the training process of PatchComplete is time-consuming. On the other hand, PatchComplete used a unified seen-category prior for all categories of data, which limited its generation ability.
@@
Continuous implicit functions represent objects as mathematical functions, such as neural network models, that implicitly describe the object’s surface as the zero level-set of the function. Chabra et al. [45] proposed to split the shape’s surface into local regions and employ independent local latent codes to represent them. To solve the imbalanced distributions between different local regions, Wang et al. [48] proposed to represent an object with clusters of local regions along with multiple decoders, where each decoder is responsible for one cluster with the same pattern.
Figure 1: Flowchart of the Prior-assisted Shape Learning Network. “Conv” and “De-Conv” denote the convolution and de-convolution operation, respectively. The PSLN leverages a Multi-scale Pattern Correlation (MPC) module to capture the associated patterns for each local region of the partial input from the priors and subsequently use these patterns to infer the complete shape hierarchically. The MPC consists of two encoders and three Cross-Attention modules. The prior encoder utilizes multiple kernels of varying sizes to capture local structures at different scales. The Cross-Attention module identifies associated local patterns by calculating attention scores between the multi-scale local features of the partial input and the priors.
@@ II-B Unsupervised 3D Shape Completion and Reconstruction
The in-domain methods [17, 18, 19, 25, 24, 23] do not use partial-complete pairs from other domains, and the data categories for the training and test set are consistent. In this paper, we mainly introduce the self-supervised point cloud methods that we have used. Mittal et al. [23] proposed the first self-supervised point cloud completion method, PointPnCNet. PointPnCNet learns to infer complete shapes from partial scans with complete ground-truth shapes as supervision. It randomly removes a region from the partial scan, employs neural networks to predict the missing part, and leverages partial matching loss and partial scans as supervision. On the one hand, the network tends to extract common patterns among shapes in the same category. On the other hand, the model is unable to distinguish between synthetic removal and natural incompleteness. As a result, PointPnCNet is capable of successfully predicting complete shapes. Wang et al. [24] proposed a two-branch cascaded refinement network for point cloud completion. In order to learn to predict complete shapes without ground truth, they remove a portion of the area from partial inputs and combine them with the MixUp [52] strategy as inputs and supervisions. Gu et al. [25] proposed a weakly-supervised model that can simultaneously predict complete shapes and 6-DoF poses with the help of multi-view partial scans. Nonetheless, the performances of these in-domain unsupervised methods are constrained as they only utilize a limited number of partial shapes from fixed categories for training.
@@ III Proposed Method
We represent the shapes as 3D voxel grids. Denote by $\mathbf{S}_{s}\in\mathbb{R}^{N\times N\times N}$ the partial shape of an object from the seen categories and $\mathbf{G}\in\mathbb{R}^{N\times N\times N}$ the corresponding ground truth, $\mathbf{S}_{u}\in\mathbb{R}^{N\times N\times N}$ the partial shape from unseen categories, where $N$ represents the resolution of the 3D voxels. Our goal is to find a mapping $f(\mathbf{S}_{u}|\mathbf{S}_{s},\mathbf{G})=\mathbf{O}$ to make the reconstructed shape $\mathbf{O}$ as close as possible to its ground truth. Our approach consists of two stages, i.e., prior-assisted Coarse Shape Learning (CoSL) and Category-specific Shape Refinement (CaSR), which will be detailed as follows.
III-A Prior-assisted Coarse Shape Learning
In real life, many 3D objects have similar local patterns. For example, most desks, benches, and lamps have sticks as support. These local structures are important for understanding the shapes of objects from unseen categories and can be used to infer missing parts of these objects. Inspired by this, we propose a Prior-assisted Shape Learning Network (PSLN) that leverages shape priors from seen categories to complete the partial scans from unseen categories.
Specifically, we first construct a prior bank ${\mathbf{B}}$ that consists of several representative complete shapes from training data. Following PatchComplete, we use the mean-shift clustering algorithm [53] to select $M$ complete shapes from training data as priors. These priors contain typical local structures from each category in the training data.
Next, we design PSLN, an end-to-end network, to learn the local patterns of the partial shape from the priors. As depicted in Figure 1, we start by using two encoders to extract local features of the partial input and priors. Both of these encoders are composed of four 3D convolutional layers. The features of the partial input and priors at the $i$-th layer are denoted by $\mathbf{X}_{i}\in\mathbb{R}^{N_{i}\times N_{i}\times N_{i}\times C_{i}}$ and $\mathbf{Y}_{i,m}\in\mathbb{R}^{N_{i}\times N_{i}\times N_{i}\times C_{i}}$, respectively, where $m=1,2,...,M$, $C$ represents the feature dimension. Then, we use cross-attention [54] to compute the weights between $\mathbf{X}_{i}$ and $\mathbf{Y}_{i,m}$ to represent their correlation. To be specific, we first split $\mathbf{X}_{i}$ and $\mathbf{Y}_{i,m}$ into vectors of size $1\times 1\times 1\times C_{i}$. We define the vectors of $\mathbf{X}_{i}$ as “queries” and the vectors of $\mathbf{Y}_{i,m}$ as “keys” and “values”. Then, we compute the dot products of each query with all keys and divide them by $C_{i}/2$. Next, we adopt a Softmax function to obtain the weights, i.e.,
$w_{i,m}=\texttt{Softmax}\left(\frac{\mathbf{X}_{i}\cdot\mathbf{Y}_{i,m}}{d/2}\right)\,.$ (1)
Finally, we assemble the prior embeddings by
@@ III-A Prior-assisted Coarse Shape Learning
where $\mathbf{V}$ represents the “values”. In practice, $\mathbf{V}$ is equal to $\mathbf{Y}$, and we only compute $\mathbf{F}_{\{2,3,4\}}$ for the features outputted by the last three encoder layers.
Additionally, we observe that local patterns of partial shapes and priors may be similar at different scales. For example, the sticks of a chair backrest may have a shape similar to that of the pole of a street lamp, but their sizes are significantly different. Although PatchComplete comprises three stages to encode multi-resolution features, it maintains the same scale for partial shapes and priors across all resolutions. As a result, it overlooks the similarity of local structures at different scales. To this end, we use convolutional kernels of various sizes to extract multi-scale features of priors and then calculate the attention between all multi-scale prior features and the features of the partial input. At the first layer of the prior encoder, we apply convolutional kernels of sizes $7\times 7\times 7$, $5\times 5\times 5$ and $3\times 3\times 3$, respectively. At the second layer, the sizes of convolutional kernels are $5\times 5\times 5$ and $3\times 3\times 3$. At the last layer, we only use a single convolutional kernel of size $3\times 3\times 3$ due to the large receptive field of this layer. The multi-scale features of each layer are concatenated and passed to the next layer. We call these blocks Multi-scale Structure Learning unit (MSL).
The assembled prior embeddings $\mathbf{F}_{i}$ are then decoded into the complete shape hierarchically. Specifically, we first directly decode $\mathbf{F}_{4}$ by a deconvolution layer. Then $\mathbf{F}_{3}$ and $\mathbf{F}_{2}$ are respectively concatenated with the decoded features and processed by two deconvolution layers. The model outputs the coarse complete shape after one more deconvolution layer and a Softmax layer. The training process of the PSLN is driven by optimizing L1 loss between the reconstructed coarse shape and its corresponding ground-truth complete shape.
Remark. Although our PSLN follows the assumption of PatchComplete, which leverages shape priors to complete the partial scans from unseen categories, the prior-utilization strategy and network architecture of our PSLN differ from PatchComplete. Specifically, PatchComplete employs a prior-learning strategy. Concretely, it trains a set of learnable shape priors and learns patch-based attention maps between these priors and partial shapes during the patch-learning stage. The target shape is generated by directly fusing the patch volumes of the learned shape priors in 3D Euclidean space. By contrast, PSLN leverages a prior-assisted strategy that uses fixed-value shape priors without requiring learning to assist in the completion of partial scans. Our goal is only to learn the multi-scale structural correlation between partial scans and priors at the feature level, then decode the correlated prior features to infer the complete shape.